OpenStudy (anonymous):

A club has 16 members. In how many ways can 3 officers consisting of a president, vice-president, and secretary be chosen? Fill in the blank with your answer which should be a positive integer

OpenStudy (anonymous):

3360

OpenStudy (anonymous):

16*15*14, assuming they can't hold more than one office. 210*16=2100+1260=3360

OpenStudy (anonymous):

why 210 times 16?

OpenStudy (anonymous):

210=15*14, so 210*16=15*14*16. I was doing it in my head

OpenStudy (anonymous):

ohhh okay so then why add 1260?

OpenStudy (anonymous):

ohhhh nevermind i see!!!!

OpenStudy (anonymous):

its a permutation question 16!/(16-3)!